OpenStudy (anonymous):

how do i find average rate of the fucntion over the gviing time. f(t)= 2t+7 [1,2]

OpenStudy (anonymous):

(f'(1)+f'(2))/2= average rate of change f'(t)= 2 (2+2)/2= 2 Also you could eye ball the question by knowing that the rate of change is constant, so the average must also be 2.